THE PHOTOELECTRON ANGULAR DISTRIBUTION AS A PROBE OF ENERGETICALLY INDISTIN-GUISABLE CHANNELS IN PHOTODETACHMENT

\underline{MATTHEW VAN DUZOR}, FOSTER MBAIWA, JIE WIE and RICHARD MABBS, Department of Chemistry, Washington University, One Brookings Dr., Campus Box 1134 Saint Louis, Missouri 63130, USA.

Photoelectron imaging measurements of I^{−}·CH_{3}CN are presented from the lowest energy I(^{2}P_{3/2})·CH_{3}CN channel threshold to 0.3 eV above the I(^{2}P_{1/2})·CH_{3}CN threshold. Excitation of the cluster just below the latter threshold leads to competition between direct detachment and the production of a dipole-bound state [I(^{2}P_{1/2})·CH_{3}CN]^{−}. Subsequent relaxation of the I moiety (^{2}P_{1/2} →^{2}P_{3/2}) occurs via autodetachment of the dipole bound electron. However, the autodetached electrons are energetically equivalent to those directly detached via the ^{2}P_{3/2} channel. While the photoelectron spectrum is insensitive to this phenomenon, the kinetic energy dependence of the photoelectron angular distribution (PAD) of the ^{2}P_{3/2} channel reflects these dynamics. Dramatic changes are observed within a 100 meV window of the I(^{2}P_{1/2})·CH_{3}CN threshold. Although previously unreported, further results are presented that show sharp changes in the PAD are relatively common in I^{−}·X cluster anions (X=polar molecule).